26th HiLumi WP2 Task Leader Meeting

Europe/Zurich
6/R-018 (CERN)

6/R-018

CERN

20
Show room on map
Videoconference Rooms
HiLumi_WP2_Task_Leader_Meeting
Name
HiLumi_WP2_Task_Leader_Meeting
Description
6th_HiLumi_WP2_Task_Leader_Meeting
Extension
109247616
Owner
Riccardo De Maria
Auto-join URL
Useful links
Phone numbers
Minutes of the 26th WP2 Task Leader Meeting held on 28/03/2014
Present: G. Arduini, O. Brüning, R. De Maria, M. Giovannozzi, P. Fessia, E. Métral, A. Valishev, E. Todesco.
 
Minutes and Follow-up of Actions (Gianluigi)
  • The minutes were approved with the following comments:
    • Massimo suggested to keep 3.5 um emittance to avoid a double standard between LHC vs HL-LHC and ion vs protons for collimation settings.  Gianluigi replied that the standard for LHC made sense because this was the physical emittance of the proton and ion beam at injection. For HL-LHC the nominal emittance for the proton beam is changed so it would make more sense to have a uniform definition at least for protons. In any case this should be decided at the level of the PLC.
    • Ezio reported that the change of geometry was the cause of the increase of peak power deposition in the triplet in the last iteration. It is not foreseen to reduce the beam clearance.
    • Ezio and Gianluigi contacted V. Baglin. The beam screen will be operated at a temperature ranging between 30 and 70 K. Elias will provide an estimate of the beam screen impedance and heat load due to image currents in the range 30 to 70 K by the end of April. Action: Elias
    • On request of Elias the deadlines for the actions on electron cloud effects (electron cloud simulations for the crab cavities and nearby transitions, and for the IP2/8 triplets, stability threshold resulting from electron clouds in the triplets/matching section in IP1/5) resulting from the Vacuum Technical meeting have been agreed for end of October so to have the results presented at the Hi-Lumi Collaboration meeting in November.
 
 
Update of the HL-LHC layout and optics – R. De Maria
  • The layout and optics aims at reaching 15 cm and 7.5/30 cm in b* with 590 mm full crossing angle in both planes and both IRs, while still keeping the magnet strengths compatible with 10 cm and 5/20cm in b* with 720 mm full crossing angle. The layout must host 4 crab cavities with 13.3 m of longitudinal space and protective masks of 1 m in from of D2, Q5, Q6 and TCTs (whose number and location needs to be confirmed) to protect the triplet, D2, Q4, Q5. The layout assumes that the cavities will be aligned within +- 0.5 mm around the nominal crossing angle orbit and beam based aligned knobs will be provided for a fine adjustment of the average orbit in the cavities. Action Stefano: clarify number and location of the TCTs.
  • The TAS aperture can be reduced by 1.5 mm in radius to reach the target of 12 sigma with imperfections or more if one wants to keep the TAS aperture only marginally larger in sigma then the triplet. Action Helmut: identify failure scenarios that require an aperture reduction of the TAS.
  • An MQY with 200 T/m at 1.9 K is compatible with the strength of the present optics in IR1 and IR5, while barely sufficient for the optics in IR6 that still needs to be validated by WP14 and WP5 for collimator settings. For IR6 adding another MQY gives more margins. IR1/5 optics with a displaced Q4 are still compatible with an MQY at 200 T/m, pending the validation of the squeeze. VDM optics at b*= 20 m are possible only at high energy. Action Riccardo: prepare the squeeze with a displaced MQY.
  • For the D2 it is proposed to maintain the aperture of 105 mm and reduce the length of the magnet (possibly to 8 m). For the beam screen it is proposed to have an octagonal beam screen with total inner horizontal and vertical clearance: 82.5 mm. The design of the beam screen and the definition of the tolerances are necessary and should be reviewed at the technical committee.
  • The orbit correctors in IR1 and IR5 must provide crossing, separation, offset in the IP, orbit correction due to triplet misalignments and transfer function errors, shift of the beam position in the crab cavity if active alignment is not provided and absorb orbit errors coming from the arcs. The required integrated field are:
    • 2.5 Tm for MCBX1-2,
    • 4.5 Tm for MCBX3,
    • 6 Tm or 7 Tm (depending on the presence/absence of active alignment of the cavities, respectively) for the correctors close to D2.
By assuming the V1.0 layout and the constraint of 3 T for orbit correctors, one would need 2.3 m long orbit correctors close to D2, while keeping orbit correctors for absorbing arc imperfections and shift beam at the crab cavities in Q4. Two other options were studied:
  • Option 2) share the strength of correctors in D2 with the one in Q4 and assuming a alignment of the cavities to the design orbit with crossing angle.
  • Option 3) install a 2.3 m vertical corrector and add a trim power converter for unbalancing D2 apertures and use D1/D2 for horizontal crossing and separation.
 
The advantage of Option 2) is to reduce length and the number the orbit corrector types in D2-Q4. The drawback is a loss of aperture in D2 TAN of about 1.5 - 1 mm. The advantage of Option 3) is to reduce the integrated lengths of the orbit correctors.
The first option cannot be implemented for space constraints unless relying on a shift of Q4 towards the arc (see later) at the cost of crab cavity voltage or aperture. In order to decide between the two remaining options we need to know:
  • Expected ramp (linear and acceleration) rate of the corrector magnets for the two options è WP3.
  • Feasibility of the active alignment of the Crab cavities required for option 2 è WP4.
  • Feasibility of unbalancing the two apertures of the  D2 magnet è WP3.
 
Gianluigi will follow-up these issues with Ezio and Rama. Action: Gianluigi
 
  • The Q4 position can be shifted towards the arc to make room for orbit correctors and crab cavities and in principle additional TCT and mask. Proofs of principle collision and injection optics have been generated with a displacement of 10 m. For its use, a smooth of the squeeze, optimization of the injection optics and VDM scan optics needs to be generated. Action Riccardo: finalize this option.
  • The movement of Q4 would allow creating space for the Long Range Wire compensator but the requirements in terms of optical functions, space need to be clarified. Action: A. Valishev, T. Pieloni.
  • The assumptions on ground motion and fiducialization errors made for the estimate of the orbit correction need to be validated by the experts. Gianluigi noted that this would be a subject for the Technical Committee.
 
Updates from Task Leaders – O. Brüning, M. Giovannozzi, E. Métral, A. Valishev
 
Task 2.2 and 2.3 (Massimo)
A squeeze sequence has been computed by Maxim: it will be finalized by Riccardo to be then discussed with WP3 in view of hysteresis effects.
Miriam progressed with the analysis of the strength requirement for the triplets and D2 correctors. A HSS presentation is planned before Easter.
Tracking studies have been performed by the SLAC colleagues at injection. A reduction of the dynamic aperture by 2 sigmas has been observed, as anticipated by Stephane, as compared to the present nominal LHC optics. A tune scan showed that it could be partially recovered. Stephane mentioned that increasing the tune split would also improve the situation. It should be possible to recover them.
Alternatives scenarios for transferring the losses resulting from Bound Free Pair Production at the IP during ion operation are being considered as an alternative to the installation of 11 T dipoles.
 
Task 2.4 (Elias)
 
 
A task 2.4 meeting has been dedicated to the longitudinal stability issues and the operation of the double harmonic RF system (800 MHz). At present an 800 MHz system is considered for:
  • Bunch flattening to reduce IBS, although the estimated impact on luminosity is marginal according to the study conducted by T. Mertens;
  • Bunch flattening in collision to minimize pile-up density (in combination with crab kissing) as proposed by S. Fartoukh;
  • Bunch flattening to reduce beam induced heating.
 
The mode of operation of the 400 MHz system in full cavity detuning mode will imply a variation of the bunch separation along the bunch train. At present no acceptable solution has been found to adjust the phase of the 800 MHz system along the bunch train accordingly. This will imply a deformation of the bunch profile that will reduce the effectiveness of the above schemes and for that reason investigations should continue to provide an operational scenario for the 800 MHz system compatible with the 400 MHz full cavity detuning mode.
A possible alternative scheme would consist in flattening the bunch profiles with band limited noise although RF noise and IBS will tend to restore Gaussian profiles.
A possible scheme using a 200 MHz Main RF system has been studied too. A 400 MHz RF system would be required as a high harmonic RF system to provide longitudinal stability. Operation in bunch lengthening mode (required to provide flat bunch profiles) would impose a limit on the bunch length to 3.4 ns to avoid loss of Landau damping.
Very likely the 200 MHz system would have to be operated in full cavity detuning mode too. Elias reminded that an ICFA mini-Workshop on “Electromagnetic wake fields and impedances in particle accelerators" will be held in Erice, from April 24th to April 28th.
Task 2.5 (Sasha)
Ji Qiang is conducting strong-strong simulations to benchmark the 2012 observations on bunch shortening. Once this is completed he should be able to start simulations to include the effect of noise of crab cavities.
Work on the possible use of an electron beam as beam-beam long range compensator is ongoing.
 
Task 2.6 (Oliver)
Nothing to report.
 
 
Next meeting will take place exceptionally on Thursday 10th April 2014 at 16:00 to avoid collision with LIU day. The status of the Beam-beam studies will be reviewed.
 
Reported by Gianluigi and Riccardo.
There are minutes attached to this event. Show them.
    • 16:00 16:20
      Approval of minutes and follow-up of actions 20m
      Speaker: Gianluigi Arduini (CERN)
    • 16:20 16:40
      Update of the HL-LHC layout and optics 20m
      Speaker: Riccardo De Maria (CERN)
      Slides
    • 16:40 17:00
      Status of the various tasks (Task leaders, 1 or 2 slides each) 20m
      Speakers: Alexander Valishev (Fermilab), Dr Elias Metral (CERN), Dr Massimo Giovannozzi (CERN), Oliver Bruning (CERN), Dr Rhodri Jones (CERN)